NicholasH:

12th grade math: Ahmed is planning to cook his family dinner and needs to purchase several food items. He is able to purchase these items either at the local grocery store or the farmer's market. Given in the table are the items and their prices. Grocery Store Farmer's Market 8 tomatoes for $13.20 10 tomatoes for $17.50 4 cups of mozzarella cheese for $4.60 3 cups of mozzarella cheese for $3.30 12 eggs for $3.00 7 eggs for $1.40 Part A: Choose one of the food items offered at the grocery store and the farmer's market and determine the unit price of the item at each location. Show all necessary work, including the name of the item chosen. (6 points) Part B: Based on your answer in part A, which location offers a better deal? Explain. (4 points)

NicholasH:

SmoothCriminal:

Part A: Let's choose the tomatoes as the victuals item for comparison. At the grocery store: 8 tomatoes for $13.20 So, the unit price at the grocery store is: \(\frac{13.20}{8} = $1.65\) per tomato. At the farmer's market: 10 tomatoes for $17.50 So, the unit price at the farmer's market is: \(\frac{17.50}{10} = $1.75\) per tomato. Part B: The grocery store offers a better deal for tomatoes because the unit price there is $1.65 per tomato, whereas the farmer's market has a unit price of $1.75 per tomato. So, the grocery store is the more cost-effective option for tomatoes.

NicholasH:

This is how you do it? you are sure? I do not want to get this wrong.

SmoothCriminal:

@nicholash wrote:
This is how you do it? you are sure? I do not want to get this wrong.
Trust me, bro. Whenever I answer a question, it's never wrong. I have to do this myself al the time. You're looking at a straight A student.

NicholasH:

Ok, well thank you man, I appreciate it.

SmoothCriminal:

@nicholash wrote:
Ok, well thank you man, I appreciate it.
Of course my guy!

SmoothCriminal:

Here's the work for Part A. Let's pick the tomatoes as the victuals item for comparison. At the grocery store: 8 tomatoes for $13.20 So, the unit price at the grocery store is: \[ \frac{13.20}{8} = $1.65 \text{ per tomato} \] At the farmer's market: 10 tomatoes for $17.50 So, the unit price at the farmer's market is: \[ \frac{17.50}{10} = $1.75 \text{ per tomato} \] Consequently, the unit price for tomatoes at the grocery store is $1.65 per tomato, and at the farmer's market, it is $1.75 per tomato.

NicholasH:

Thank you! :)

SmoothCriminal:

No problemo mi hermano